Documenta Mathematica

Extra Volume Quadratic Forms and Related Topics, LSU, Baton Rouge, 2001

Preface  1
List of Talks  3
List of Participants  7

Jón Kr. Arason

Witt Groups of Projective Line Bundles  11–48

Ricardo Baeza Some Algebraic Aspects of Quadratic Forms over Fields

of Characteristic Two  49–63

Karim Johannes Becher On the Number of Square Classes

of a Field of Finite Level  65–84

V. Chernousov, V. Guletskiǐ 2-Torsion of the Brauer Group of an Elliptic Curve:

Generators and Relations  85–120

A. Dress, K. T. Huber, V. Moulton Metric Spaces

in Pure and Applied Mathematics  121–139

Robert W. Fitzgerald

Isotropy and Factorization in Reduced Witt Rings  141–163

Alexander Hahn The Zassenhaus Decomposition for the Orthogonal Group:

Properties and Applications  165–181

Detlev W. Hoffmann Dimensions of Anisotropic

Indefinite Quadratic Forms, I  183–200

Max-Albert Knus and Oliver Villa Quadratic Quaternion Forms,

Involutions and Triality  201–218

Ahmed Laghribi Certaines Combinaisons Linéaires de Deux Formes de Pfister

et le Problème d’Isotropie  219–240

David W. Lewis, Claus Scheiderer, Thomas Unger A Weak Hasse Principle for Central Simple

Algebras with an Involution  241–251

iv